# Leveraging LLMs for Enhanced Open-Vocabulary 3D Scene Understanding in Autonomous Driving

### v1

@@ 4.1 Dataset
We use a subset of the WayveScenes101 dataset [28], designed specifically to advance novel view synthesis in challenging driving scenarios. We selected a subset of five representative scenes (IDs: 12, 21, 36, 78 and 88) totaling about 500 images. This selection balances computational feasibility with the need for diverse and challenging data, encompassing various environmental conditions, lighting variations, and complex urban environments. This allows us to thoroughly evaluate our approach in handling the complexities of autonomous driving environments. For quantitative evaluation of our method’s semantic understanding capabilities, we generated segmentation maps for the images in the evaluation set using a pre-trained Segformer model [29] on the Cityscapes dataset [30]. This semantic information allows us to assess semantic accuracy.
@@ 4.2 Implementation
We trained the LE3DGS using the same loss function as described in [21] on the Wayvesene101 dataset. We set $\lambda_{s}=0.5$, $\lambda_{CE}=\lambda_{smo}=\lambda_{u}=0.1$ and other weights to the deafault. The training process involved two main phases: dense semantic feature extraction and 3D Gaussian point optimization. For the semantic feature extraction, we set the number of epochs to 500, ensuring a rich semantic representation of the scene. In the optimization phase for 3D Gaussian points, we employed a 2D position gradient threshold of 0.0001, with a total of 20,000 iterations. The densification process was halted halfway through, at 10,000 iterations. All other hyperparameters were kept consistent with those used in [2]. The entire training process was conducted on a single RTX 4090 GPU, taking approximately 40 minutes to complete for each scene in the dataset. The mean peak signal-to-noise ratio (PSNR) achieved is around 25.14.
Table 1: Input structure for GPT-3.5 Turbo query generation in our implementation. This setup guides the language model to provide structured information about objects relevant to driving scenarios.
Role Content
System A helpful assistant for a computer vision task, providing structured information about objects to pay attention to while driving. Include both important objects and nearby objects that might be at the borders. Instructions for formatting the response, including: Main Positive (object to pay attention to) 1-4 Helping Positives (related terms or attributes) 4-6 Negatives (objects to differentiate from, including similar objects, nearby objects, and background elements)
User "Driving through an intersection in an {Road Type} on a {Weather} {Time of Day}, what objects should the driver pay attention to? " or show the {Object}. "
@@
For the inference phase, we leveraged the API of GPT-3.5 Turbo [31] to dynamically generate queries. For each primary query, the model also provided 1 to 4 helping positive words and 4 to 6 canonical phrases. Table 1 denotes the input structure to the LLM model. The {Road Type}, {Weather}, and {Time of Day} in the user prompts could be provided from the dataset scene metadata [28]. We then calculated the relevancy score using Algorithm 1. To obtain object segmentations, we applied a threshold of 0.5 to the computed relevancy scores. Figure 1 illustrates our inference approach pipeline for scene 88 in Wayvescene101 dataset. The canonical phrases used in the baseline approach are “object”, “things”, “stuff”, and “texture”, consistent with the approach described in [21].
@@
Ours w/o $p_{\texttt{help}}$
Ours
@@
4.3 Results
Our approach demonstrates significant improvements in open-vocabulary segmentation across various urban driving scenarios. Figure 2 provides a visual comparison, showcasing the progression from overly broad segmentation using predefined canonical phrases to refined results with our full LLM-enhanced method. Our approach accurately identifies and segments various elements such as trees, traffic lights, and sidewalks with high precision, illustrating the benefits of LLM-generated context-aware queries and helping positives.
Inference Method $\mathrm{Accuracy}\uparrow$ $\mathrm{Presicion}\uparrow$ $\mathrm{mIoU}\uparrow$ $\mathrm{mAP}\uparrow$
Predefined Canonical Phrase 0.65 0.33 0.23 0.49
Ours w/o $p_{\texttt{help}}$ 0.84 0.42 0.33 0.53
Ours 0.92 0.58 0.37 0.54
Table 2: Comparison of inference methods: our approach (with and without helping positive words $p_{\texttt{help}}$) versus predefined canonical phrases.
Quantitative results in Table 2 support these visual observations. Our full method achieves the highest performance across all metrics, with notable improvements in Accuracy (0.92), Precision (0.58), and mean Intersection over Union (mIoU) (0.37). These represent significant gains over the predefined canonical phrase method (0.65, 0.33, and 0.23 respectively). The ablation study (Ours w/o $p_{\texttt{help}}$) indicates the robustness of our LLM-driven query generation, outperforming the baseline even without helping positive words.
(a) Mean Accuracy
(b) Mean IoU
(c) Mean Precision
Figure 3: Performance comparison of our methods against baseline across different object classes for the scenes in dataset.
Figure 3 further breaks down performance across various urban scene elements. Our method shows significant improvements in Mean IoU and Mean Precision for challenging classes like "sidewalk", "traffic light", and "traffic signs", while maintaining an edge in common classes such as "building" and "car". These results collectively demonstrate the effectiveness of our LLM-enhanced approach in improving scene understanding and object segmentation across diverse urban environments.
For more detailed qualitative results, please refer to section A.1 in the appendix.
5 Discussion
Our approach demonstrates significant improvements in open-vocabulary 3D scene understanding, but there are several important aspects to consider:
Limitations of 3D Gaussian Splatting
While 3D Gaussian Splatting is effective for certain scenes, it struggles with generating large and dynamic scenes [32]. This limitation can affect the overall performance and applicability of our approach in more dynamic environments, such as busy urban intersections or highways with fast-moving vehicles.
System Integration and Real-time Performance
The integration of our approach with existing autonomous driving systems presents both opportunities and challenges. While our method offers enhanced semantic understanding, its real-time performance in the context of rapidly changing driving scenarios requires further optimization. Current autonomous systems often rely on faster, albeit less semantically rich, perception methods. Bridging this gap involves not only improving our method’s computational efficiency but also developing strategies for seamless integration with existing sensor fusion and decision-making pipelines.
@@ 6 Conclusion
In this study, we proposed a novel method that integrates LLMs with LE3DGS to improve autonomous vehicle scene understanding. By leveraging the robust reasoning capabilities of LLMs and the detailed spatial representation of 3D Gaussians, our approach significantly improves the accuracy and precision of open vocabulary object segmentation and scene interpretation tasks. Our experimental results indicate that our method outperforms traditional predefined canonical phrase-based approaches on all aspects of accuracy, precision, and mean Intersection over Union.
Future work will focus on enhancing our system’s capability to handle dynamic scenes by integrating LLMs with advanced 3D Gaussian Splatting techniques. We aim to fuse our approach with multi-modal sensor data, such as LiDAR, to create a more robust and comprehensive perception system. Another crucial next step is to seamlessly integrate our advanced semantic understanding with autonomous driving planners. This integration will involve developing efficient interfaces that translate our rich semantic insights into actionable information, directly informing path planning and decision-making algorithms for safer and more intelligent autonomous navigation.
Appendix A Appendix
A.1 Qualitative Results
Figures 4, 5, and 6 showcase our approach’s effectiveness in open vocabulary segmentation across diverse driving scenes. Each figure presents, from top to bottom: ground truth images, Segformer-generated segmentation maps (used for quantitative evaluation), our high-quality scene reconstructions using LE3DGS, and our open vocabulary segmentation results. Our method accurately highlights key elements like cars, pedestrians, trees, buildings, and traffic signs/lights. For quantitative evaluation, we align with Segformer classes [33], but our visualizations include additional classes to demonstrate the flexibility of our approach. These figures illustrate our method’s capability to accurately reconstruct and semantically segment complex urban scenes, even for classes beyond traditional segmentation models.
Figure 4: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scene 88.
(a) Scene 12
(b) Scene 21
Figure 5: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scenes 12 and 21.
(a) Scene 36
(b) Scene 78
Figure 6: Comparison of ground truth, segmentation by Segformer, reconstruction, and our open vocabulary segmentation results (top to bottom) for scenes 36 and 78.

### v3

@@ 4.1 Dataset
4.1.1 WayveScene
We use a subset of the WayveScenes101 dataset [46], designed specifically to advance novel view synthesis in challenging real-world driving scenarios. We selected a subset of five representative scenes (IDs: 12, 21, 36, 78 and 88) totaling about 500 images. This selection balances computational feasibility with the need for diverse and challenging data, encompassing various environmental conditions, lighting variations, and complex urban environments. This allows us to thoroughly evaluate our approach in handling the complexities of autonomous driving environments. For quantitative evaluation of our method’s semantic understanding capabilities, we generate segmentation maps for the images in the evaluation set using a pre-trained Segformer model [35] on the Cityscapes dataset [6]. This semantic information allows us to assess the semantic accuracy of our proposed method.
4.1.2 Text Dataset
We generate a set of helping positives and canonical phrases for each query object using GPT-3.5 Turbo [25] for fine-tuning. For each primary query, the model provided 1 to 3 helping positive words and up to 6 canonical phrases. Figure 3 denotes the structure of one example for scene 12 from WayveScene101 in the generated dataset. The “Scene” description could be provided by the user or using an image captioning model.
4.2 LLMs
In addition to GPT-3.5 Turbo, we employ a variety of smaller, open-source models to improve inference efficiency and support on-device computation for applications such as autonomous vehicles and robotics. Specifically, we utilize instruction-tuned Qwen2.5 [37, 29], available in parameter sizes of 0.5, 1.5, 3, and 7 billion, along with instruction-tuned Llama 3.2 [10], offered in 1 and 3 billion parameter configurations. To optimize performance further, we fine-tune these models on a custom text dataset that we generated from GPT-3.5 Turbo.
4.2.1 Fine-tuning Process
We fine-tune the instruction-tuned Qwen and Llama models using LoRA [13] (Low-Rank Adaptation) for efficient parameter updates. We leverage Unsloth [7], which optimizes LoRA’s backward propagation, reducing FLOPs during gradient descent. The key projection layers (q_proj, k_proj, v_proj, o_proj) were targeted with a rank of 16 and a scaling factor of 16. The training is performed with a batch size of 2, gradient accumulation of 4 steps, and a learning rate of $1\times 10^{-4}$ using the AdamW optimizer and capped at 25 epochs. Mixed precision (FP16/BF16) accelerates training, while regular evaluations and early stopping ensure convergence. This optimized approach efficiently trains the models for generating canonical phrases and helping positives.
(a) Qwen 0.5B Instruct (top) and fine-tuned (bottom)
(b) Qwen 1.5B Instruct (top) and fine-tuned (bottom)
(c) Qwen 3B Instruct (top) and fine-tuned (bottom)
(d) Qwen 7B Instruct (top), fine-tuned (bottom)
(e) Llama 1B (top) and fine-tuned (bottom)
(f) Llama 3B (top) and fine-tuned (bottom)
Figure 4: Qualitative comparison of semantic segmentation results before and after fine-tuning for different model variants. Each subfigure shows three views of the same scene, with the top row representing results from the instruction-tuned model (before fine-tuning) and the bottom row showing results from the fine-tuned model.
@@ 4.3 Implementation
We train the LE3DGS using the same loss function as described in [28] on the Wayvesene101 dataset. We set $\lambda_{s}=0.5$, $\lambda_{CE}=\lambda_{smo}=\lambda_{u}=0.1$ and other weights to the default. The training process involved two main phases: dense semantic feature extraction and 3D Gaussian point optimization. For the semantic feature extraction, we set the number of epochs to 500, ensuring a rich semantic representation of the scene. In the optimization phase for 3D Gaussian points, we employed a 2D position gradient threshold of 0.0001, with a total of 20,000 iterations. The densification process was halted halfway through, at 10,000 iterations. All other hyperparameters were kept consistent with those used in [19]. The entire training process was conducted on a single RTX 4090 GPU, taking approximately 40 minutes to complete for each scene in the dataset. The mean peak signal-to-noise ratio (PSNR) achieved is around 25.14.
@@
For the inference phase, we leveraged the API of GPT-3.5 Turbo [25] and the small Instruct-tuned and fine-tuned Qwen and Llama models running on the device to generate queries, helping positives and canonical words dynamically. We then calculated the relevancy score using Algorithm 1. We applied a threshold of 0.5 to the computed relevancy scores to obtain object segmentations. The canonical phrases used in the baseline approach are “object”, “things”, “stuff”, and “texture”, consistent with the approach described in [28].
4.4 Results
We evaluated the performance of Qwen and Llama model variants and GPT-3.5 Turbo against the baseline approach for open vocabulary segmentation through Intersection over Union (IoU), Accuracy, Precision, and mean Average Precision (mAP) metrics. Our experimental results, presented in Table 1, demonstrate that GPT-3.5 Turbo significantly outperforms the baseline approach across all metrics, achieving an IoU of $0.27$, accuracy of $0.90$, and precision of $0.37$ without helping positives, compared to the baseline’s IoU of $0.18$, accuracy of $0.64$, and precision of $0.19$. This substantial improvement validates the effectiveness of leveraging large language models for open vocabulary segmentation tasks.
The fine-tuned variants of both Qwen and Llama models show marked improvements over their instruction-tuned counterparts. Notably, the Qwen2.5-1.5B-Finetuned model achieves the highest IoU ($0.28$) and precision ($0.37$) among all models in the configuration without helping positives, matching GPT-3.5 Turbo’s performance. Similarly, the Llama3.2-3B-Finetuned model demonstrates strong performance with an IoU of $0.26$ and accuracy of $0.84$, suggesting that effective fine-tuning can elevate the performance of smaller models to the levels approaching that of much larger models. Figure 4 shows the qualitative comparison of fine-tuned and instruct-tuned models.
In our ablation study examining the quality of generated helping positives ($P_{help-pos}$), we observe that a model’s capacity significantly affects its ability to generate effective helping words. Larger models generate more semantically meaningful helping positives, leading to stable or improved segmentation performance, while smaller variants generate less effective helping words that can degrade segmentation quality. For instance, as shown in Table 1, the helping positives generated by Qwen2.5-7B-Finetuned enable robust segmentation performance (IoU: $0.23$, Accuracy: $0.89$, Precision: $0.34$), while helping positives from smaller models lead to more significant performance drops in the segmentation pipeline.
@@ 4.4 Results
GPT without $p_{\texttt{help-pos}}$
GPT with $p_{\texttt{help-pos}}$
@@
The quality of helping positives reaches its peak with GPT-3.5 Turbo, whose generated helping words enable substantial improvements in segmentation performance, achieving IoU ($0.32$, $+18.5\%$), accuracy ($0.94$, $+4.4\%$), and precision ($0.50$, $+35.1\%$). This marked enhancement suggests that larger-scale models can generate more contextually appropriate and semantically rich helping positives, leading to more effective scene segmentation. The quantitative results are further supported by qualitative comparisons shown in Figure 5, which demonstrate improved segmentation quality across various urban scene elements.
@@ 5 Conclusion
In this study, we proposed a novel method that integrates LLMs with LE3DGS to enhance scene understanding for autonomous vehicles. Our approach significantly improves open vocabulary object segmentation in complex driving environments, substantially outperforming traditional predefined canonical phrase-based approaches. Our results demonstrate that carefully fine-tuned smaller language models can achieve performance comparable to larger expert models while enabling on-device applications. Furthermore, our ablation studies revealed that the effectiveness of helping positives correlates with model scale, with larger models better equipped to leverage additional semantic information. Future work will focus on integrating this semantic understanding with autonomous driving planners, developing efficient interfaces that translate semantic insights into actionable information for path planning and decision-making algorithms.